Suppose the price of a stock is $50 at the beginning of the year, the risk-free rate is 2%, assumed constant and $2 dividend is to be paid after half a year. For a long forward position with delivery in one year, find its value after 9 months if at that time the stock price turns out to be

a) $52

b) $55
Answer #1

Answer:

Given,

Forward value after 9 months = 50*(1+rf)^T + 2*(1+rf)^(T)

= 50*(1 + 0.02)^(9/12) + 2*(1 + 0.02)^(3/12)

= 50*1.02^(9/12) + 2*1.02^(3/12)

= 50.75 + 2.01

= 52.76

a)

If price = 52

Value = 52 - 52.76

= - 0.76

b)

If price = 55

Value = 55 - 52.76

= 2.24